Solve the differential equation by an appropriate method. y''+2y'=x^2+4e^2x


1
Expert's answer
2021-04-15T07:19:41-0400

We have given the differential equation,


"y''+2y'=x^2+4e^{2x}"


"(D^2+2D)y = x^2 + 4e^{2x}"

Auxiliary equation : "m^2 + 2m = 0"

"m(m+2) = 0"

"m = 0, m = -2"

Hence, "C.F = C_1 + C_2e^{-2x}"


"P.I = \dfrac{1}{D^2+2D}4e^{2x} + \dfrac{1}{D^2+2D}x^2"


"= \dfrac{4e^{2x}}{2^2+2(2)} + \dfrac{1}{2D}[1+\dfrac{D}{2}]^{-1}x^2"


"= \dfrac{4e^{2x}}{8} + \dfrac{1}{2D}[1-\dfrac{D}{2}+\dfrac{D^2}{4}]x^2"


"= \dfrac{4e^{2x}}{8} + [\dfrac{1}{2D}-\dfrac{1}{4}+\dfrac{D}{8}]x^2"


"P.I = \dfrac{e^{2x}}{2} + \dfrac{x^3}{6} - \dfrac{x^2}{4} + \dfrac{x}{4}"


Hence, complete solution of the differential equation is


"y = C.F + P.I"


"y = C_1 + C_2e^{2x} + \dfrac{e^{2x}}{2} + \dfrac{x^3}{6} - \dfrac{x^2}{4} + \dfrac{x}{4}"